Mathematics 241 – Multivariable Calculus Quiz 2 – September 13, 2013

Your Name:

A) (10) Find parametric equations for the line through the point (2, 2, 1) in the direction of (1, 0, −2).

B) (10) Find the equation of the plane in R^{3} containing p = (1, 1, 0) and orthogonal to the line {(1, 1, 0) + t(3, 4, −1) | t ∈ R}.